Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
Paoletti, FIori, Tringali - 21:05 Thursday 08 April 2021 (51355) Print this report
Finding a noisy device in CEB

The recent IPS ENEL outage occured Saturday the 3rd of April forced some systems to shut-down. Automatically, of manually for safety reasons.

As visible in the plot (right down image, the three currents of the CEB UPS line), there were current "spikes" in phase "S" before the IPS ENEL outage.

These spikes are of a device that activates every two hours, and only for 19 seconds. Everyday. And it is powered by one of the three UPS phases only.
It sinks a current of about 3Arms, therefore is a device of about 700VA

This device is potentially "dangerous", because when activated it creates many sidebands in the UPS power line, and these are seen also in the CEB magnetometers.
From past experience, we know that having this type of noise could be a problem in Hrec (we have seen this with many other devices).

This periodic noise has come back three days later, Tuesday morning, around 9:30 LOCAL TIME (maintenance period). Probably with a manual action done for recovering from an automatic switch off.

Have someone did something at that time? Has any devices (maybe some chiller) been turned ON again just before that time?
Please report to the ENV group any action.

Images attached to this report
Comments to this report:
tacca - 18:25 Thursday 08 April 2021 (51356) Print this report

I would change the last sentence in:
Please, report ON THE LOGBOOK any action.

dattilo - 10:41 Friday 09 April 2021 (51367) Print this report
Triggered by Federico, I made some investigation, and in the end it turns out that that the mysterious device is... the CEB's Clean Rooms AHU.
I asked Davide, and he told me that he switched ON the AHU at 08:57 LT of Tuesday April 6th. (it remained OFF at the time of the ENEL problem, and Davide forgot to switch it ON the morning after...).
In the attached plot the effect of the power ON is clearly visible.
Images attached to this comment
dattilo - 13:37 Saturday 10 April 2021 (51380) Print this report
It is quite strange that the Clean Rooms AHU, which is powered by the IPS, produces effects on the UPS line. So, to be sure it is really the AHU, I tried to reproduce the event.
In agrement with the Control Room and Davide, I then turned OFF the AHU on 09-04 at 14:24 UTC, and turned it back ON today at 08:50, by turning ON the same general switch (labeled E1) on which Davide had operated.
From the attached plots it can be seen that glitches are also present during the shutdown period.
Therefore, the device that is producing the glitches is something else.
It is worth mentioning that in the period 09:03 - 09:33 UTC the UPS_CURR_S was lower, as the lights of the L-shaped corridor adjacent to the DET area were OFF.
Images attached to this comment
Search Help